OpenStudy (anonymous):

A smoker is 15 times more likely to die of lung cancer than a non-smoker. An ex smoker who stopped smoking t years ago is w times more likely to die of lung cancer than a non-smoker where... w=15-t How much more likely is Sandy to die of lung cancer than polly who quite 2 1/2 years ago?

OpenStudy (ganpat):

given equation , w=15-t and t = 2.5 can you calculate w ???

OpenStudy (ganpat):

??

OpenStudy (anonymous):

I am having trouble figuring out sandy's

OpenStudy (anonymous):

or over thinking it idk w=12.5 but that's polly's risk not sandy's right?

OpenStudy (ganpat):

so is sandy smoker or not ?? its not mentioned, i guess, he is..

OpenStudy (anonymous):

sandy quit 2 1/2 years ago

OpenStudy (anonymous):

polly has never smoked

OpenStudy (ganpat):

u sure ?? it seems, polly quit 2.5 yrs ago

OpenStudy (anonymous):

Graph the equation and use the graph to estimate how much more likely it is for Sandy to die of lung cancer than Polly, if Polly never smoked and Sandy quit yr ago.

OpenStudy (anonymous):

copied and paseted : ) lol Im going to paste the whole problem next

OpenStudy (anonymous):

A smoker is 15 times more likely to die of lung cancer than a nonsmoker. An ex-smoker who stopped smoking t years ago is w times more likely to die of lung cancer than a nonsmoker, where • w = 15 – t. Graph the equation and use the graph to estimate how much more likely it is for Sandy to die of lung cancer than Polly, if Polly never smoked and Sandy quit yr ago.

OpenStudy (ganpat):

h ah a, ok.. polly never smoked, so sandy should die 12.5 yrs earlier than polly.. right ??

OpenStudy (anonymous):

so polly is origin

OpenStudy (ganpat):

yup...

OpenStudy (anonymous):

0,0 ?

OpenStudy (anonymous):

how do I graph 12.5? is it x or y im still lost

OpenStudy (ganpat):

now graph.. W = 15 - t... t=0, w = 15 t=5, w = 10 t = 10, w = 5..  does that makes sense ?/